OpenStudy (anonymous):

solve log(2) (8-6x)=5

OpenStudy (anonymous):

x= -4

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Change into exponential form: 2^5=8-6x -6x=24 x= -4